A car is traveling at 9 miles per second. The car's speed increases to 18 miles per second in 3 seconds as it passes a truck ahead of it. What is the acceleration of the car?
Show your work.

Answer by Alan3354(69443)

A car is traveling at 9 miles per second. The car's speed increases to 18 miles per second in 3 seconds as it passes a truck ahead of it. What is the acceleration of the car?
----------
Accel = (v1 - v0)/t
Accel = (18-9)/3 = 3 miles/sec/sec
==============
That's a fast car with admirable acceleration.
The car does 18 mi/sec = 64,800 mi/hr, and maybe could go faster.
That would put NASCAR out of business.
------------
I would pay big bucks for that truck, even.